OpenStudy (anonymous):

does anyone know what geometric progressions and arithmetic progressions are?

OpenStudy (anonymous):

An arithmetic progression is a list of numbers where the dierence between successive numbers is constant. The terms in an arithmetic progression are usually denoted as u1; u2; u3 etc. where u1 is the initial term in the progression, u2 is the second term, and so on; un is the nth term. AND A geometric progression is a list of terms as in an arithmetic progression but in this case the ratio of successive terms is a constant. In other words, each term is a constant times the term that immediately precedes it.

OpenStudy (anonymous):

can you give me an example

OpenStudy (anonymous):

An example of an arithmetic progression is 2; 4; 6; 8; 10; 12; 14; : : : Since the dierence between successive terms is constant, we have u3 - u2 = u2 - u1 and in general un+1 - un = u2 - u1 We will denote the dierence u2 - u1 as d, which is a common notation.

OpenStudy (anonymous):

An example of a geometric progression is Find r for the geometric progression whose rst three terms are 2, 4,8